Required information [The following information applies to the questions displayed below.) Summary information from the finan1A Currer 1A Invent Turnover Ratio Acct Rec Turn 1A Invent Turnover 1A Days Sal in Inv 1A Days Sal Uncol 1B short term For bo1A Current Ratio 1A Acid Test Ratio 1A Acct Rec Turn 1A Invent Turnover 1A Days Sal in Inv 1A Days Sal Uncol 1B short term Fo1A Current Ratio 1A Acid Test Ratio 1A Acct Rec Turn 1A Invent Turnover 1A Days Sal in Inv 1A Days Sal Uncol 1B short term Fo1A Days Sal Uncol 1B short term 00 1A Current 1A Acid Test 1A Acct Rec 1A Invent 1A Days Sal in Ratio Ratio Turn Turnover Inv1A Current Ratio 1A Acid Test Ratio 1A Acct Rec Turn 1A Invent Turnover 1A Days Sal in Inv 1A Days Sal Uncol 1B short term Fo1A Current Ratio 1A Acid Test Ratio 1A Acct Rec Turn 1A Invent Turnover 1A Days Sal in 1A Days Sal Uncol 1B short term Inv Fo1A Acid Test Ratio 1A Current Ratio 1A Acct Rec Turn | 1A Invent Turnover 1A Days Sal in 1A Days Sal Uncol 1B short term Iny
Answer #1
1A Current ratio= Current assets / Current liability
Barco $148,940 / $71,340               2.09
Kyan $231,650 / $91,300               2.54
Acid Test ratio = (Current assets - Prepaid expenses - merchandise inventory) / Current liability
(Cash + Account receivable) / Current liability
Barco (18000+40400) / $71,340
Barco 58400 / $71,340               0.82
Kyan (35000+54400) / $91,300
Kyan 89400 / $91,300               0.98
Accounts receivables turnover Net Annual Credit Sales ((Beginning Accounts Receivable + Ending Accounts Receivable) / 2)
Barco 800,000 / ((29,800+40,400)/2)            22.79 Times
Kyan 882,200 / (57200+54400)/2            15.81 Times
Inventory turnover Cost of Goods Sold / ((Beginning Inventory + Ending Inventory) / 2)
Barco 595,100 / (53,600+85,040)/2 8.58
Kyan 648,500 / (111,400+134,500)/2               5.27
Computation of days sales inventory Inventory / COGS x 365 = Days sales inventory
Barco 85040 / 595100 x 365 =       52.16
Kyan 134500 / 648500 x 365 =       75.70
Days sales uncollectable
Account receivable / Net sales x 365 = Days sales uncollected
Barco 40400 / 800000 x 365 =       18.43
Kyan 54400 / 882200 x 365 =       22.51
Kyan has better short term credit risk it's current ratio and acid test ratio is higher.